### v1

# MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models

Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across multi-modal tasks by scaling model size and training data. However, these dense LVLMs incur significant computational costs and motivate the exploration of sparse Mixture of Experts (MoE) architectures. While MoE improve parameter efficiency, effectively applying MoE to simultaneously model modality-specific features and cross-modal associations in LVLMs remains challenging. In this work, we propose to incorporate Mixture of Intra- and Inter-Modality Experts (MoIIE) to LVLMs. For each token, expert routing is guided by its modality, directing tokens to their respective intra-modality experts as well as a shared pool of inter-modality experts, enabling the model to jointly learn rich intra-modal features and cross-modal interactions. We further introduce an effective and straightforward two-stage training strategy, which facilitates the direct activation of both MoE and multi-modal capabilities. Extensive experiments across different data scales and LLM backbone demonstrate the effectiveness, efficiency and generality of our approach. Notably, our MoIIE models with 5.5B and 11.3B activated parameters match or even surpass the performance of existing advanced open-source MoE-LLMs based multi-modal models that involve more activated parameters. The code is available at https://github.com/AlenjandroWang/MoIIE.
1 Introduction
Large Vision-Language Models (LVLMs) Bai et al. (2023b); Dai et al. (2023); Liu et al. (2023a); Zhu et al. (2023) have gained significant attention for their ability to process information across both visual and linguistic modalities Cui et al. (2023); Li et al. (2024d). By integrating visual encoders Radford et al. (2021); Zhai et al. (2023a) with Large Language Models (LLMs) through connection module Lin et al. (2024b), LVLMs align high-dimensional visual features with the linguistic knowledge and reasoning capabilities of LLMs Bai et al. (2023a); Chiang et al. (2023); Bi et al. (2024), demonstrating effectiveness across diverse cross-modal tasks Liu et al. (2024c); Fu et al. (2024).
As with unimodal LLMs, scaling up model size has been shown to improve performance in multi-modal settings Liu et al. (2024b); Chen et al. (2024c) but also significantly increases computational costs, especially when using dense Transformer Vaswani et al. (2023) architectures. To maintain efficiency while scaling parameters, recent studies introduce Mixture of Experts (MoE) Lepikhin et al. (2020) into LLMs, replacing dense feed-forward network (FFN) layers with sparsely activated expert layers. This approach adaptively selects only a small subset of experts for each input based on token-level routing decisions, thus reducing computational overhead while enhancing model capacity.
For multi-modal MoE implementation, a common approach is to directly extend vanilla MoE designs from LLMs to LVLMs by routing tokens from all modalities to a shared pool of experts Lin et al. (2024a); Han et al. (2025). However, this overlooks the fundamental differences in information density and feature distribution between text and image tokens Liang et al. (2023). An alternative approach incorporates modality-specific experts, where text and image tokens are routed to their respective specialized expert groups. While this design enables more specialized feature learning for each modality Lin et al. (2024c); Wang et al. (2024b); Shen et al. (2023), these experts primarily focus on intra-modal knowledge and neglect cross-modal associations, such as the alignment between noun tokens in text and corresponding entity regions in images Xiao et al. (2024), as illustrated in Figure 1 (a).
To this end, we proposes the Mixture of Intra- and Inter-Modality Experts (MoIIE), to simultaneously capture both modality-specific features and cross-modal associations in LVLMs. As illustrated in Figure 1 (b), MoIIE comprises three distinct groups of experts: two intra-modality groups that specialize in language and vision, independently processing text and image tokens; and an inter-modality expert group shared by both modalities that focuses on cross-modal interactions between text and image tokens. Correspondingly, MoIIE learns two dedicated routers respectively for image and text tokens, with each token dynamically routed to the most relevant experts from the corresponding intra-modality group and the inter-modality group. As not all tokens are strongly associated across modalities, such as function words in text or visual regions without descriptive content, the routing mechanism allows for flexible combinations. Some tokens may activate only intra-modality experts or only inter-modality experts, while others may access both expert groups simultaneously. This dynamic routing strategy is consistently applied to both image and text tokens.
Additionally, for multi-modal MoE training, most existing approaches adopt a sparse upcycling strategy Komatsuzaki et al. (2023) that transforms dense LVLMs into sparse models. This strategy generally follows a three-stage training pipeline: multi-modal pre-training of the dense model, fine-tuning the dense model on diverse multitasking datasets with all parameters unfrozen, and model sparsification with only expert modules trained Lin et al. (2024a); Chen et al. (2024b). However, this pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose that multi-modal fine-tuning and sparsification can be jointly optimized with all parameters updated in a more straightforward and effective two-stage training strategy. In the first stage, visual inputs are aligned with the LLM backbone through multi-modal pre-training datasets. In the second stage, MoIIE is integrated and the entire model is optimized using fine-tuning datasets, to simultaneously learn (1) versatile multi-modal understanding and instruction-following capabilities, while enabling (2) expert learning of modality-specific knowledge with cross-modal association within the MoIIE module. Our experimental results demonstrate that this simplified strategy is easy to implement and also effectively enhances performance across diverse downstream tasks.
Figure 1: Modality-specific MoE v.s. Our MoIIE. Left: The modality-specific MoE module routes text and image tokens exclusively to their respective specialized expert groups, limiting cross-modal associations such as the alignment between “dog” token and its corresponding image region. Right: Our MoIIE introduces intra-modality and inter-modality expert groups. Intra-modality experts (Expert V for image tokens, Expert L for text tokens) process modality-specific features while inter-modality experts (Expert S) process tokens from both modalities to model cross-modal interactions.
In summary, our contributions are threefold:
  • We propose a robust sparse LVLM framework equipped with the novel MoIIE module, effectively modeling both modality-specific features and cross-modal associations.
  • We introduce an effective and straightforward two-stage training strategy that simultaneously optimizes multi-modal fine-tuning and MoE modules.
  • We conduct extensive experiments demonstrating that our MoIIE-integrated model through the two-stage training strategy achieves superior scaling efficiency compared to existing dense, modality-expert-based, and original MoE-based LVLMs.
Figure 2: Method overview. Left: The MoIIE architecture, consisting of intra-modality experts (Expert V for image tokens and Expert L for text tokens) and inter-modality experts (Expert S) that process tokens from both modalities. Right: The two-stage training strategy. For each module, the tunable or frozen icon before the slash indicates the configuration during the first stage, while the icon after the slash represents the second-stage setup.
2 Methodology
2.1 Overview
We introduce MoIIE, a multi-modal MoE variant integrated into LVLMs to capture both modality-specific features and cross-modal associations, with the comprehensive architecture outlined in Figure 2. It incorporates a pretrained visual encoder coupled with a connection module that transforms visual inputs into sequential representations matching the dimension of LLM token embeddings. A detailed explanation of the MoIIE architecture is provided in Section 2.2, followed by our proposed simple two-stage training paradigm in Section 2.3. A detailed comparison between the original MoE and our MoIIE is shown in Appendix, highlighting the key architectural innovations improving multi-modal learning.
2.2 MoIIE Architecture
2.2.1 Multi-modal Input Representation
We process and represent inputs from different modalities into sequential embeddings compatible with LLMs. Specifically, given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the original resolution, we use a pre-trained visual encoder followed by an MLP-based connection module to extract and project image features into the LLM embedding space as: $X_{I}=\text{MLP}(\text{Encoder}(I))$, resulting in a sequence representation of image tokens $X_{I}=[x_{1}^{\mathcal{I}},\dots,x_{m}^{\mathcal{I}}]\in\mathbb{R}^{m\times d}$, where $m=h\times w$ represents the number of image tokens. The MLP module ensures dimensional alignment with the LLMs’ embedding space for seamless integration. For textual inputs $T\in\mathbb{Z}^{L}$, we tokenize and embed them to obtain a sequence of textual embeddings $X_{T}=[x_{1}^{\mathcal{T}},\dots,x_{n}^{\mathcal{T}}]\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of text tokens. The final multi-modal input representation $X$ is obtained by concatenating the visual and textual embeddings: $X=[X_{I},X_{T}]\in\mathbb{R}^{(m+n)\times d}$, which is then fed into the LLM for joint processing of multi-modal information.
2.2.2 Sparse MoIIE Forward
The MoIIE is integrated into the LLM to facilitate efficient and adaptive multi-modal learning. The expert modules comprise intra-modality and inter-modality expert groups, formally defined as $\mathcal{E}=[E_{1}^{\mathcal{T}},\dots,E_{L}^{\mathcal{T}},E_{L+1}^{\mathcal{I}},\dots,E_{3L}^{\mathcal{I}},E_{3L+1}^{\mathcal{S}},\dots,E_{4L}^{\mathcal{S}}]$, where intra-modality expert groups $E^{\mathcal{T}}$ and $E^{\mathcal{I}}$ respectively process text and image tokens for learning modality-specific features. In contrast, the inter-modality expert group $E^{\mathcal{S}}$ processes tokens from both modalities to capture rich cross-modal associations. To balance modality-specific and shared representations, we establish the following expert allocation constraints: $|E^{\mathcal{T}}|+|E^{\mathcal{I}}|=|E^{\mathcal{S}}|=2L$ and $|E^{\mathcal{T}}|=|E^{\mathcal{I}}|=L$. This configuration ensures that the number of intra-modality experts equals the number of inter-modality experts, while maintaining equal allocation to each modality’s experts.
The token routing mechanism is governed by modality-specific routers $G^{\mathcal{M}}(x)$, which include a visual router $G^{\mathcal{I}}(x)$ and a textual router $G^{\mathcal{T}}(x)$. These routers dynamically assign each input token to its top-$K$ most relevant experts based on predicted activation probabilities as follows:
$G^{\mathcal{M}}(x_{n})=\text{Softmax}(\text{top-}K(x_{n}\cdot W_{g}^{\mathcal{M}})),$ (1)
[…]
$\text{MoIIE}(x_{n})=\begin{cases}\sum_{i=1}^{K}G_{i}^{\mathcal{T}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{T}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{T}\\ \sum_{i=1}^{K}G_{i}^{\mathcal{I}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{I}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{I}\end{cases}$ (2)
where each token $x_{n}$ is processed by its top-$K$ activated experts $\text{E}_{i}(x_{n})$, with their outputs aggregated through a probability-weighted summation. Specifically, image tokens are routed by the visual router $G^{\mathcal{I}}(x)$ to the top-$K$ experts selected from the combined pool of intra-vision and inter-modality groups $E^{\mathcal{I/S}}$, while textual tokens are routed by the textual router $G^{\mathcal{T}}(x)$ to the top-$K$ experts from the intra-language and inter-modality groups $E^{\mathcal{T/S}}$.
The routing mechanism enables flexible expert selection patterns: each token may activate only intra-modality experts, only inter-modality experts, or a combination of both. This hierarchical routing strategy ensures efficient multi-modal learning while simultaneously modeling both modality-specific features and cross-modal associations. All experts in $\mathcal{E}$ are initialized from the LLM’s FFN layers, preserving prior knowledge while enabling adaptive multi-modal specialization.
2.3 Training Recipe
To balance simplicity and efficacy, we propose a two-stage training strategy that effectively activates the capabilities of the MoIIE architecture. In the first stage, we focus on pretraining the connection module to align visual representations with the LLM’s linguistic embedding space, enabling consistent processing of inputs across different modalities. In the second stage, we initialize both intra-modality and inter-modality experts using the pre-trained FFNs from the base LLM and fine-tune the entire model with diverse visual instruction datasets. During this stage, the training objectives include multi-modal expert learning for capturing modality-specific features and cross-modal associations, and optimizing versatile multi-modal understanding and instruction-following capability.
Training Objective
To ensure load balancing among experts within the MoIIE module while optimizing the model for overall multi-modal understanding, the training objective $\mathcal{L}$ is composed of two components: the language modeling cross-entropy loss $\mathcal{L}_{\text{lm}}$ and an auxiliary loss $\mathcal{L}_{\text{aux}}$ aimed at load balancing among experts MistralAITeam (2023).
$\mathcal{L}=\mathcal{L}_{\text{lm}}+\alpha\cdot\mathcal{L}_{\text{aux}},$ (3)
where $\alpha$ is a weighting coefficient set to 0.001. Notably, higher values of $\alpha$ adversely impact model performance, as detailed in Section 3.6. The auxiliary loss $\mathcal{L}_{\text{aux}}$ is defined as:
$\mathcal{L}_{\text{aux}}=|\mathcal{E}|\cdot\sum_{i=1}^{|\mathcal{E}|}\left(\mathbb{E}_{x}\left[G_{i}(x)\right]\cdot\mathbb{E}_{x}\left[\mathbf{1}_{i}(x)\right]\right),$ (4)
$G_{i}(x)$ is the routing probability of assigning token $x$ to expert $E_{i}$, and $\mathbf{1}_{i}(x)$ denotes the indicator function that equals 1 when expert $E_{i}$ is activated for token $x$ and 0 otherwise.
3 Experiments
3.1 Implementation Details
We employ a pre-trained SigLIP model Zhai et al. (2023b) as the visual encoder and a two-layer MLP as the connection module. For the LLM backbone, we utilize phi-3-mini Marah Abdin (2024) and LLaMA3-8b AI@Meta (2024). Our method and all compared MoE architectures use a 4-experts configuration. Specifically, our MoIIE module includes 2 intra-modality experts (one for vision and one for language) and 2 inter-modality experts. More detailed discussion of expert settings is provided in the ablation studies 3.6.
[…]
Data Backbone General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
Dense 2M Phi-3-mini 74.4 62.7 81.6 42.9 70.2 86.2 28.3 65.1 41.9 53.2 64.3 41.3 30.8 57.1
Vanilla MoE 2M Phi-3-mini 75.1 62.8 81.5 42.8 70.6 86.3 28.8 65.1 42.2 53.1 65.2 40.6 29.9 57.2
Modality MoE 2M Phi-3-mini 75.1 62.9 81.7 41.0 71.0 86.2 29.3 66.0 42.7 53.1 65.5 40.4 31.0 57.4
MoIIE 2M Phi-3-mini 75.3 63.2 81.8 42.8 71.2 86.6 30.5 65.4 42.9 53.3 65.9 41.3 31.9 57.9
With More Visual Instruction Data
Dense 2.7M Phi-3-mini 75.0 63.4 81.6 41.1 70.1 87.0 31.7 64.6 48.2 57.5 73.7 40.1 31.0 58.8
Vanilla MoE 2.7M Phi-3-mini 75.2 63.3 81.4 42.2 70.6 87.2 30.8 65.2 47.8 57.2 74.7 42.2 31.2 59.1
Modality MoE 2.7M Phi-3-mini 73.9 63.5 81.7 40.1 70.4 87.1 32.9 65.7 48.9 58.8 74.1 40.4 31.3 59.1
MoIIE 2.7M Phi-3-mini 75.9 64.1 82.0 42.7 71.8 87.2 30.7 66.9 48.5 58.8 75.9 42.4 32.5 60.0
With Larger LLM Backbone
Dense 2M LLaMA3-8B 74.8 64.8 82.2 47.9 72.1 86.6 30.1 67.2 43.4 54.0 76.6 41.1 30.9 59.4
Vanilla MoE 2M LLaMA3-8B 75.5 64.5 82.1 46.0 71.6 86.3 33.0 67.3 43.8 54.7 76.2 40.7 30.6 59.4
Modality MoE 2M LLaMA3-8B 75.6 64.3 82.3 46.3 72.3 86.7 35.0 67.6 44.9 55.4 76.7 41.6 31.2 60.0
MoIIE 2M LLaMA3-8B 75.7 64.9 82.3 47.5 73.0 87.0 36.5 67.9 44.7 56.0 76.9 42.8 32.2 60.6
During training, we utilize the AdamW optimizer Kingma & Ba (2017) with a cosine learning rate scheduler for one epoch across both stages. In the first stage, only the connection module is optimized with a learning rate of $1\times 10^{-3}$, using the Bunny-pretrain-LAION-2M dataset He et al. (2024). In the second stage, all parameters are unfrozen for joint SFT and sparse upcycling training. The learning rate is set to $2\times 10^{-6}$ for the visual encoder and $2\times 10^{-5}$ for all other components. For this stage, we attempt training with different scales of visual instruction data: (1) 1.3M samples from MGM-Instruction Li et al. (2024c); (2) 2M samples by adding Bunny-695K He et al. (2024); (3) 2.7M samples by further incorporating LLaVA-NEXT-779k Liu et al. (2024b). Training is conducted using DeepSpeed Shuaiwen Leon Song (2023) with ZeRO-3 optimization. All compared baselines in the main results and our method follow the same training setup and data configurations. Further implementation details are provided in Appendix A.
3.2 Evaluation Benchmarks
We conduct a comprehensive evaluation across 13 diverse multi-modal benchmarks. Specifically, general multi-modal benchmarks include MMBench-EN Liu et al. (2024c), MM-Vet Yu et al. (2024), GQA Hudson & Manning (2019), VQAv2 and SEED-Image  Li et al. (2023a). For knowledge-based question answering, we utilize MMMU validation subset Yue et al. (2024), AI2D Kembhavi et al. (2016),SciQA-IMG Lu et al. (2022) and Mathvista Lu et al. (2024b). For OCR-based question answering, we assess performance on TextVQA Singh et al. (2019), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021). Additionally, we evaluate hallucination robustness using POPE Li et al. (2023c) and HallusionBench Guan et al. (2024).
3.3 Main Results
We compare our MoIIE-equipped model against three baseline variants: the original dense model (Dense) without MoE, the vanilla MoE model MistralAITeam (2023) that routes tokens from all modalities to a shared pool of experts (Vanilla MoE), and the model with modality-specific experts where text and image tokens are routed to their respective specialized experts (Modality MoE). The comparisons are conducted across multiple visual instruction tuning data scales and varying LLM backbones, as shown in Table 1. The key findings are as follows:
(1) The MoIIE module consistently outperforms all other architectural variants across most benchmarks with different data scales and LLM backbone sizes, demonstrating that integrating both modality-specific features and cross-modal associations can effectively enhance performance on diverse multi-modal tasks. Further analysis of expert modules is provided in Appendix A.
(2) Specifically, our MoIIE module is particularly effective on knowledge-based QA and hallucination benchmarks that requiring sophisticated cross-modal interaction between textual concepts and corresponding visual entity regions. In contrast, OCR-based QA tasks primarily involve interpreting information verbatim from images, where Modality MoE can already achieves competitive performance. This performance difference highlights MoIIE’s capacity to handle both modality-specific features and cross-modal associations.
(3) We further provide detailed performance trends with increasing fine-tuning data in Figure 3. As the data scale continues to increase, our MoIIE progressively improves while other architectures, especially Dense and Modality MoE, encounter performance limitations. Moreover, with larger training datasets, our MoIIE consistently outperforms all alternatives, with the performance gap widening between MoIIE and both Dense and Modality MoE. This suggests that the MoIIE framework offers superior scaling properties for multi-modal learning, effectively leveraging larger datasets to enhance representation power without the parameter inefficiency of dense models or the limited cross-modal reasoning of strictly modality-separated experts.
Figure 3: Performance variation across different visual instruction tuning data scales . MoIIE outperforms other architectural variants in achieving superior scaling efficiency.
3.4 Comparison with MoE-based LLMs Initialized LVLMs
We further compare our MoIIE pre-trained on the LLaVA-OV-Single-Image dataset Li et al. (2024a) with existing open-sourced LVLMs initialized from MoE-based LLMs to demonstrate its effectiveness. While SPHINX-X Liu et al. (2024a), MGM Li et al. (2024c) and CuMO Li et al. (2024b) that are built on Mixtral-8x7B MistralAITeam (2023) leverage large-scale language data to provide robust initialization and enhanced language capabilities, they typically demand significant computational resources, rendering them less practical for resource-constrained settings. In contrast, our approach achieves comparable performance on diverse multi-modal benchmarks with 17% fewer activated parameters, reducing overall parameter count and lowering deployment and training costs. Furthermore, our approach allows for the construction of MoE-based LVLMs from any LLM backbones, supporting more flexible expert settings without sacrificing performance.
Table 2: Comparisons of MoIIE with open-source LVLMs utilizing pre-trained MoE-based LLM backbones. In addition to performance gains, our MoIIE offers the flexibility to be constructed using any dense LLM backbone. “Act. Param” indicates activated parameters while “All Param” indicates the total number of model parameters.
[…]
MoE-based LVLMs model with pre-trained MoE LLM backbone
SPHINX-X Liu et al. (2024a) - $\sim$40B 68.0 63.8 81.1 89.6 71.3 40.9 31.1 42.7 74.5
MGM Li et al. (2024c) 13.5B $\sim$40B 69.2 - - - 75.6 45.8 41.8 41.8 -
CuMO Li et al. (2024b) 13.5B $\sim$40B 66.0 63.8 81.8 85.7 75.3 48.7 45.0 38.2 77.9
MoE-based LVLMs model from dense LLM backbone
MoIIE-A5B 5.5B 7B 69.5 63.4 82.5 87.1 77.8 42.6 43.6 46.4 90.1
MoIIE-A11B 11.3B 16B 70.5 64.5 82.9 87.3 77.2 47.3 42.7 46.0 89.4
3.5 The Discussion of Training Pipeline
Converting dense LLMs into MoE-based LVLMs typically involves a three-stage training process Lin et al. (2024a); Chen et al. (2024b). First, multi-modal pre-training that transforms dense LLMs into dense LVLMs by aligning visual and textual representations. The second stage is to fine-tune the dense LVLMs with multi-modal instruction data followed by a multi-modal sparse upcycling phase to obtain MoE-based LVLMs. However, this three-stage training pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose a streamlined two-stage training strategy that consolidates SFT and model sparsification into a single stage after pretraining, enabling the direct activation of both multi-modal and MoE capabilities.
The three stages training variants utilize 2M visual instruction data for SFT with all parameters tuned, and LLaVA-NEXT-779k for sparse upcycling with only MoE layers tuned. The two stages training variants utilize 2.7M visual instruction data for joint SFT and sparse upcycling to tune all model parameters. As shown in Table 3, our proposed two-stage training method consistently outperforms the previous three-stage training variants under the same training settings across comprehensive multi-modal benchmarks, demonstrating that our approach is simpler and also more effective.
Table 3: Comparisons of different training strategies. Pretrain-SFT-MoE represents the three-stage training method where MoE denotes sparse upcycling from multi-modal dense checkpoints, tuning only MoE layers Lin et al. (2024a) after SFT. Pretrain-SFT&MoE is our proposed two-stage training method, in which sparse upcycling is integrated with SFT, tuning the full model parameters simultaneously.
[…]
Method Stage General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
Pretrain-SFT-MoE 3 75.6 63.1 81.6 41.2 70.3 86.9 31.0 64.8 48.2 56.2 63.8 39.3 31.3 57.9
Pretrain-SFT&MoE 2 75.7 63.5 81.8 42.7 71.2 87.2 32.1 65.3 49.6 58.8 65.3 41.9 32.5 59.0
3.6 Ablation Study
Table 4: Ablation study for various MoIIE module configurations. This table compares the performance of different MoIIE setups, including variations in the MoIIE location, load balance coefficient, number of experts and experts settings.
[…]
Number of Experts 4 8 65.1 75.6 63.4 42.2 86.5 43.0 71.2 63.9(+0.2)
Experts Balance(8 experts) Balance Unbalance 65.4 75.0 63.4 41.1 86.8 41.3 71.6 63.5(-0.2)
MoIIE Module Location Interleaved Full 65.8 75.8 63.1 40.3 87.0 42.6 70.6 63.6(-0.1)
Load Balance Coefficient 0.001 0.01 65.5 75.4 63.2 39.0 86.4 42.0 70.7 63.2(-0.5)
Ours(MoIIE-A5B) - - 65.4 75.3 63.2 41.3 86.6 42.8 71.1 63.7
The Impact of the Number of Experts
Increasing the number of experts typically improves MoE performance  Lepikhin et al. (2020); FedusF et al. (2022). In Table 4, we vary the total number of experts in the MoIIE module while keeping the number of activated parameters and the allocation ratio of different expert groups constant. Results show that expanding from 4 to 8 experts results in performance gains of 0.3, 0.2, and 0.9 points on MMBench, GQA, MM-Vet, and MMMU, respectively, while performance on POPE and TextVQA remains stable or slightly decreases. These results suggest that a larger expert pool can improve generalization on complex benchmarks, but may require more extensive data to fully leverage its capacity to achieve performance improvement.
The Impact of Experts Balance
Maintaining a balanced allocation of intra-modality and inter-modality experts proves beneficial for performance. As shown in Table 4, we examine the impact of expert balance under an 8-expert configuration. In the Balanced setting, we maintain a 1:1 ratio between intra-modality and inter-modality experts, specifically consisting of 2 vision experts, 2 language experts, and 4 shared experts. Each router selects the top-2 experts from its corresponding intra-modality pool and the inter-modality expert pool, resulting in 2 + 4 candidate experts. In contrast, the unbalanced setting adopts a 3:1 ratio, with 3 vision experts, 3 language experts, and only 2 shared experts. In this configuration, each router selects the top-2 experts from 3 intra-modality and 2 inter-modality candidates. Experimental results indicate that the balanced setting consistently outperforms the unbalanced configuration, suggesting that maintaining an even distribution of intra- and inter- modality experts yields more effective cross-modal interaction and overall performance.
Layer Location of MoIIE module
We explore two architectural configurations for integrating the MoIIE module into LVLMs: the “interleaved” configuration and the “full” configuration following the designs in Lin et al. (2024a); AI@Meta (2025). In the interleaved setup, MoIIE module are incorporated every other layer, interspersed with standard dense layers throughout the model, maintaining a balance between sparse and dense computations. In contrast, the full configuration replaces all layers with MoIIE layers, maximizing the use of sparse experts across the entire network.
As shown in Table 4, fully replacing all layers with MoIIE modules does not necessarily lead to significant performance improvements but substantially increases training costs and computational demands due to the larger parameter count. In contrast, the interleaved configuration achieves comparable performance while maintaining higher training efficiency and reduced resource consumption. Therefore, we adopt the interleaved configuration as the optimal design choice for MoIIE-based LVLMs, ensuring a well-balanced trade-off between performance, efficiency, and scalability.
The Impact of Load Balance Coefficient
In MoE architectures, there is an inherent trade-off between load balance and model performance. The auxiliary loss is designed to promote load balance among experts Cai et al. (2024). However, we observe that while enforcing balance, this loss inevitably interferes with multi-modal learning, potentially hindering the model’s capacity to capture cross-modal associations. As shown in Table 4, we evaluate MoIIE under different values of the auxiliary loss weight $\alpha$. The results indicate that increasing $\alpha$ leads to a noticeable decline in overall performance, particularly on MMMUval and MM-Vet, where the drop is most pronounced. This suggests that finding an optimal balance is essential to maintain robust multi-modal understanding, ensuring that expert specialization is not excessively constrained by strict load balancing.
3.7 Visualization of Token Pathways
Figure 4: Visualization of experts activated pathways. The figure shows the top-2 activated experts for text and image tokens, with Expert V and Expert L are intra-modality experts, Expert S denotes inter-modality experts.
As illustrated in Figure 4, we visualize the expert activation pathways for image and text tokens during the feed-forward process in the MoIIE module on the MME test set. Our analysis reveals distinct activation patterns across different model layers. In the shallow layers (the first half), both image and text tokens predominantly activate intra-modality experts in their respective modalities (Expert V for vision and Expert L for text). We hypothesize that this behavior is due to shallow layers focusing on extracting modality-specific features with minimal cross-modal interaction.
In contrast, the deeper layers of the model (the second half) show a shift toward activating inter-modality experts (Expert S), promoting deeper cross-modal interactions. Interestingly, both image and text tokens tend to activate the same inter-modality experts in these deep layers. For example, in the 9th MoIIE layer, Expert S_1 is frequently activated, while in the 12th MoIIE layer, Expert S_2 becomes the dominant expert. These findings indicate that our MoIIE module effectively balances modality-specific and cross-modal processing, dynamically adjusting expert activation to facilitate both early-stage feature extraction and late-stage modality fusion.
4 Related Works
Large Vision-Language Models
Recent advancements in large language models (LLMs)  Bai et al. (2023a); AI@Meta (2024); Touvron et al. (2023b); Chiang et al. (2023); Jiang et al. (2023); Bi et al. (2024); Team (2023); OpenAI (2023b, a); Touvron et al. (2023a) have demonstrated remarkable generalization and instruction-following capabilities, driving progress toward general-purpose AI. To extend these abilities to both linguistic and visual domains, research on Large Vision-Language Models (LVLMs) has gained increasing attention Li et al. (2023b); Alayrac et al. (2022); Liu et al. (2023a, b, 2024b). Recent efforts in LVLMs have improved performance through high-quality data He et al. (2024); Chen et al. (2024a); Zhang et al. (2024), additional training phases Bai et al. (2023b); Lu et al. (2024a); Laurençon et al. (2024), high-resolution input support Li et al. (2024e); Wang et al. (2024b), and multi-encoder enhancements Shi et al. (2024); Liu et al. (2024a); Fan et al. (2024). The latest open-source LVLMs Li et al. (2024a); Wang et al. (2024a); Chen et al. (2025) leverage dynamic resolution, large-scale datasets, and advanced LLMs to achieve state-of-the-art results.
Mixture of Experts in LLMs
The MoE paradigm Jacobs et al. (1991) enables selective expert activation, allowing for efficient scaling while maintaining manageable computational costs. Integrating MoE into LLMs involves replacing feed-forward networks with expert modules, using a Top-K router for activation Lepikhin et al. (2020); Du et al. (2022); FedusF et al. (2022); Zoph et al. (2022); Rajbhandari et al. (2022); Xue et al. (2024). Sparse upcycling Komatsuzaki et al. (2023) reduces costs by transforming dense models into sparse ones. Recent MoE models xAI (2024); MistralAITeam (2023); Muennighoff et al. (2024); DeepSeek-AI (2024a, b) refine performance through large-scale data and enhanced training strategies.
Mixture of Experts in LVLMs
MoE is used in LVLMs, such as ARIA Li et al. (2025) and DeepSeek-VL2 Wu et al. (2024), leveraging pre-trained MoE-based LLM backbones MistralAITeam (2023); DeepSeek-AI (2024a) but are limited by fixed expert settings and routing mechanisms along with high training costs. Sparse upcycling as in  Chen et al. (2024b); Lin et al. (2024a, c) are utilized to improve scalability but still faces challenges in capturing both modality-specific features and cross-modal relationships. Zhou et al. (2025) introduce extend the traditional Parameter-efficient fine-tuning framework to support multi-modal expert, but the impact of the modality expert groups has not been thoroughly explored. To address these issues, we propose a modality-aware MoE architecture, which models both intra-modality experts and inter-modal experts. Our two-stage training strategy optimizes MoE activation, offering a more flexible and scalable approach for building high-performance LVLMs.
5 Conclusion
In this study, we propose MoIIE, a novel Mixture-of-Experts architecture paired with a simple yet effective two-stage training strategy that enables the flexible construction of powerful MoE-based LVLMs from any dense LLM. MoIIE organizes experts into specialized intra-modality and shared inter-modality groups to create dedicated pathways for both visual and linguistic processing while enabling rich cross-modal associations, significantly enhancing performance across a wide range of multi-modal tasks. Extensive experiments demonstrate that our framework achieves superior scaling efficiency compared to existing dense and sparse LVLM architectures.
Broader Impacts and Limitations
A primary limitation of our work lies in the relatively limited training data and its exclusive focus on vision-language modalities. In future work, we plan to address these limitations by incorporating more high-quality multi-task datasets, supporting dynamic input resolutions, and extending MoIIE to broader modalities such as speech, thereby advancing its generalization and applicability in real-world multi-modal scenarios.
Appendix A Appendix
Figure 5: Comparison between the Vanilla MoE and our MoIIE. Left: The vanilla MoE module MistralAITeam (2023) routes all modality tokens into a single group of experts. Right: The MoIIE module introduces intra-modality and inter-modality experts group. Intra-modality Expert V for image tokens and Expert L for text tokens modeling modality specfic features. Inter-modality experts (Expert S) that process tokens from both modalities, modeling cross-modal associations, capable of handling tokens from both modalities. The visual router routes image tokens to Expert V and Expert S, while the textual router routes text tokens to Expert L and Expert S.
Configuration Stage 1 Stage 2
Experts type - FFN
Experts -
1 visual expert
2 shared experts
1 textual expert
Top-K - 2
Visual encoder siglip-so400m-patch14-384
Connection module 2 Linear layers with GeLU
Image resolution 384 x 384
Learning rate
1e-3
{connection module}
2e-5
{LLM,connection module}
2e-6
{visual encoder}
LR schedule Cosine decay
Weight decay 0
Optimizer AdamW
Warmup ratio 0.03
Epoch 1
Global batch size 256 128
Deepspeed Zero2 Zero3
Max token length 4096
Table 5: Detailed training hyperparameters of MoIIE.
A.1 Visualization of activated experts
As illustrated in the figure6, we visualize the proportion of activated experts across four representative multimodal understanding benchmarks. For general QA, we use MME Fu et al. (2024); for OCR-based QA, TextVQA Singh et al. (2019); for visual-centric reasoning, we use MMVP Tong et al. (2024); and for hallucination, we use HallusionBench Guan et al. (2024).
We consistently observe that lower layers of the model tend to activate modality-specific visual experts, indicating specialized modeling of image features. In contrast, as the layer depth increases, the activation ratio of modality-specific experts decreases, while the use of modality-shared experts increases, suggesting a shift toward modeling shared semantic information across modalities.
Figure 6: Visualization of the proportion of activated experts across diverse dimensions in multimodal understanding benchmarks.
A.2 Training Details
Our detailed training settings and hyper-parameters of MoIIE are shown in Table 5.
A.3 Expert Analysis
MMBench TextVQA MMMU_val
Textual experts 72.9 63.1 41.5
Visual experts 73.6 64 41
Shared experts 74.1 63.1 41.3
Table 6: Expert analysis between modality-specific experts and modalitt-shared experts.
To evaluate the effectiveness of modality-specific and modality-shared experts, we extracted checkpoints from the shared experts in MoIIE, as well as the modality-specific visual and textual experts from the Modality Expert model. These were tested on the MMbench Liu et al. (2024c) and TextVQA Singh et al. (2019) and MMMU validation subset Yue et al. (2024) which for General, OCR and konwledge based abilities of model.
As shown in Table 6 We find that modality-shared experts outperform modality-specific experts on general multi-modal benchmarks like MMbench Liu et al. (2024c), which require cross-modal reasoning. In contrast, visual experts perform better on OCR-based QA like TextVQA Singh et al. (2019), which emphasizes image details and modality-specific information whether texual experts are capable of knowledge-based QA like MMMU validation subset Yue et al. (2024), which requires the original textual knowledge of model. This further validates the effectiveness of our MoIIE architecture, which demonstrates that successful performance across a wide range of multi-modal tasks requires modeling both cross-modal associations and modality-specific features.