### v2

# MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models

Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across multi-modal tasks by scaling model size and training data. However, these dense LVLMs incur significant computational costs and motivate the exploration of sparse Mixture of Experts (MoE) architectures. While MoE improve parameter efficiency, effectively applying MoE to simultaneously model modality-specific features and cross-modal associations in LVLMs remains challenging. In this work, we propose to incorporate Mixture of Intra- and Inter-Modality Experts (MoIIE) to LVLMs. For each token, expert routing is guided by its modality, directing tokens to their respective intra-modality experts as well as a shared pool of inter-modality experts, enabling the model to jointly learn rich intra-modal features and cross-modal interactions. We further introduce an effective and straightforward two-stage training strategy, which facilitates the direct activation of both MoE and multi-modal capabilities. Extensive experiments across different data scales, visual encoder and LLM backbone demonstrate the effectiveness, efficiency and generality of our approach. Notably, our MoIIE models with 5.5B and 11.3B activated parameters match or even surpass the performance of existing advanced open-source MoE-LLMs based multi-modal models that involve more activated parameters.
MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models
Dianyi Wang1,2  Siyuan Wang3  Zejun Li1  Yikun Wang1,2 Yitong Li4  Duyu Tang4  Xiaoyu Shen5  Xuanjing Huang1  Zhongyu Wei1,2 1Fudan University  2Shanghai Innovation Institut 3University of Southern California  4Huawei Technologies Co., Ltd 5Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT dywang24@m.fudan.edu.cn, sw_641@usc.edu
1 Introduction
Large Vision-Language Models (LVLMs) Bai et al. (2023b); Dai et al. (2023); Liu et al. (2023a); Zhu et al. (2023) have gained significant attention for their ability to process information across both visual and linguistic modalities Cui et al. (2023); Li et al. (2024d). By integrating visual encoders Radford et al. (2021); Zhai et al. (2023a) with Large Language Models (LLMs) through connection module Lin et al. (2024b), LVLMs align high-dimensional visual features with the linguistic knowledge and reasoning capabilities of LLMs Bai et al. (2023a); Chiang et al. (2023); Bi et al. (2024), demonstrating effectiveness across diverse cross-modal tasks Liu et al. (2024c); Fu et al. (2024).
As with unimodal LLMs, scaling up model size has been shown to improve performance in multi-modal settings Liu et al. (2024b); Chen et al. (2024c) but also significantly increases computational costs, especially when using dense Transformer Vaswani et al. (2023) architectures. To maintain efficiency while scaling parameters, recent studies introduce Mixture of Experts (MoE) Lepikhin et al. (2020) into LLMs, replacing dense feed-forward network (FFN) layers with sparsely activated expert layers. This approach adaptively selects only a small subset of experts for each input based on token-level routing decisions, thus reducing computational overhead while enhancing model capacity.
For multi-modal MoE implementation, a common approach is to directly extend vanilla MoE designs from LLMs to LVLMs by routing tokens from all modalities to a shared pool of experts Lin et al. (2024a); Han et al. (2025). However, this overlooks the fundamental differences in information density and feature distribution between text and image tokens Liang et al. (2023). An alternative approach incorporates modality-specific experts, where text and image tokens are routed to their respective specialized expert groups. While this design enables more specialized feature learning for each modality Lin et al. (2024c); Wang et al. (2024b); Shen et al. (2023), these experts primarily focus on intra-modal knowledge and neglect cross-modal associations, such as the alignment between noun tokens in text and corresponding entity regions in images Xiao et al. (2024), as illustrated in Figure 1 (a).
To this end, we proposes the Mixture of Intra- and Inter-Modality Experts (MoIIE), to simultaneously capture both modality-specific features and cross-modal associations in LVLMs. As illustrated in Figure 1 (b), MoIIE comprises three distinct groups of experts: two intra-modality groups that specialize in language and vision, independently processing text and image tokens; and an inter-modality expert group shared by both modalities that focuses on cross-modal interactions between text and image tokens. Correspondingly, MoIIE learns two dedicated routers respectively for image and text tokens, with each token dynamically routed to the most relevant experts from the corresponding intra-modality group and the inter-modality group. As not all tokens are strongly associated across modalities, such as function words in text or visual regions without descriptive content, the routing mechanism allows for flexible combinations. Some tokens may activate only intra-modality experts or only inter-modality experts, while others may access both expert groups simultaneously. This dynamic routing strategy is consistently applied to both image and text tokens. Additionally, for multi-modal MoE training, most existing approaches adopt a sparse upcycling strategy Komatsuzaki et al. (2023) that transforms dense LVLMs into sparse models. This strategy generally follows a three-stage training pipeline: multi-modal pre-training of the dense model, fine-tuning the dense model on diverse multitasking datasets with all parameters unfrozen, and model sparsification with only expert modules trained Lin et al. (2024a); Chen et al. (2024b). However, this pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose that multi-modal fine-tuning and sparsification can be jointly optimized with all parameters updated in a more straightforward and effective two-stage training strategy. In the first stage, visual inputs are aligned with the LLM backbone through multi-modal pre-training datasets. In the second stage, MoIIE is integrated and the entire model is optimized using fine-tuning datasets, to simultaneously learn (1) versatile multi-modal understanding and instruction-following capabilities, while enabling (2) expert learning of modality-specific knowledge with cross-modal association within the MoIIE module. Our experimental results demonstrate that this simplified strategy is easy to implement and also effectively enhances performance across diverse downstream tasks.
Figure 1: Modality-specific MoE v.s. Our MoIIE. Left: The modality-specific MoE module routes text and image tokens exclusively to their respective specialized expert groups, limiting cross-modal associations such as the alignment between “dog” token and its corresponding image region. Right: Our MoIIE introduces intra-modality and inter-modality expert groups. Intra-modality experts (Expert V for image tokens, Expert L for text tokens) process modality-specific features while inter-modality experts (Expert S) process tokens from both modalities to model cross-modal interactions.
In summary, our contributions are threefold:
  • We propose a robust sparse LVLM framework equipped with the novel MoIIE module, effectively modeling both modality-specific features and cross-modal associations.
  • We introduce an effective and straightforward two-stage training strategy that simultaneously optimizes multi-modal fine-tuning and MoE modules.
  • We conduct extensive experiments demonstrating that our MoIIE-integrated model through the two-stage training strategy achieves superior scaling efficiency compared to existing dense, modality-expert-based, and original MoE-based LVLMs.
Figure 2: Method overview. Left: The MoIIE architecture, consisting of intra-modality experts (Expert V for image tokens and Expert L for text tokens) and inter-modality experts (Expert S) that process tokens from both modalities. Right: The two-stage training strategy. For each module, the tunable or frozen icon before the slash indicates the configuration during the first stage, while the icon after the slash represents the second-stage setup.
2 Methodology
2.1 Overview
We introduce MoIIE, a multi-modal MoE variant integrated into LVLMs to capture both modality-specific features and cross-modal associations, with the comprehensive architecture outlined in Figure 2. It incorporates a pretrained visual encoder coupled with a connection module that transforms visual inputs into sequential representations matching the dimension of LLM token embeddings. A detailed explanation of the MoIIE architecture is provided in Section 2.2, followed by our proposed simple two-stage training paradigm in Section 2.3. A detailed comparison between the original MoE and our MoIIE is shown in Appendix, highlighting the key architectural innovations improving multi-modal learning.
2.2 MoIIE Architecture
2.2.1 Multi-modal Input Representation
We process and represent inputs from different modalities into sequential embeddings compatible with LLMs. Specifically, given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the original resolution, we use a pre-trained visual encoder followed by an MLP-based connection module to extract and project image features into the LLM embedding space as: $X_{I}=\text{MLP}(\text{Encoder}(I))$, resulting in a sequence representation of image tokens $X_{I}=[x_{1}^{\mathcal{I}},\dots,x_{m}^{\mathcal{I}}]\in\mathbb{R}^{m\times d}$, where $m=h\times w$ represents the number of image tokens. The MLP module ensures dimensional alignment with the LLMs’ embedding space for seamless integration. For textual inputs $T\in\mathbb{Z}^{L}$, we tokenize and embed them to obtain a sequence of textual embeddings $X_{T}=[x_{1}^{\mathcal{T}},\dots,x_{n}^{\mathcal{T}}]\in\mathbb{R}^{n\times d}$, where $n$ denotes the number of text tokens. The final multi-modal input representation $X$ is obtained by concatenating the visual and textual embeddings: $X=[X_{I},X_{T}]\in\mathbb{R}^{(m+n)\times d}$, which is then fed into the LLM for joint processing of multi-modal information.
2.2.2 Sparse MoIIE Forward
The MoIIE is integrated into the LLM to facilitate efficient and adaptive multi-modal learning. The expert modules comprise intra-modality and inter-modality expert groups, formally defined as $\mathcal{E}=[E_{1}^{\mathcal{T}},\dots,E_{L}^{\mathcal{T}},E_{L+1}^{\mathcal{I}},\dots,E_{3L}^{\mathcal{I}},E_{3L+1}^{\mathcal{S}},\dots,E_{4L}^{\mathcal{S}}]$, where intra-modality expert groups $E^{\mathcal{T}}$ and $E^{\mathcal{I}}$ respectively process text and image tokens for learning modality-specific features. In contrast, the inter-modality expert group $E^{\mathcal{S}}$ processes tokens from both modalities to capture rich cross-modal associations. To balance modality-specific and shared representations, we establish the following expert allocation constraints: $|E^{\mathcal{T}}|+|E^{\mathcal{I}}|=|E^{\mathcal{S}}|=2L$ and $|E^{\mathcal{T}}|=|E^{\mathcal{I}}|=L$. This configuration ensures that the number of intra-modality experts equals the number of inter-modality experts, while maintaining equal allocation to each modality’s experts.
The token routing mechanism is governed by modality-specific routers $G^{\mathcal{M}}(x)$, which include a visual router $G^{\mathcal{I}}(x)$ and a textual router $G^{\mathcal{T}}(x)$. These routers dynamically assign each input token to its top-$K$ most relevant experts based on predicted activation probabilities as follows:
$G^{\mathcal{M}}(x_{n})=\text{Softmax}(\text{top-}K(x_{n}\cdot W_{g}^{\mathcal{M}})),$ (1)
[…]
$\text{MoIIE}(x_{n})=\begin{cases}\sum_{i=1}^{K}G_{i}^{\mathcal{T}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{T}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{T}\\ \sum_{i=1}^{K}G_{i}^{\mathcal{I}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{I}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{I}\end{cases}$
where each token $x_{n}$ is processed by its top-$K$ activated experts $\text{E}_{i}(x_{n})$, with their outputs aggregated through a probability-weighted summation. Specifically, image tokens are routed by the visual router $G^{\mathcal{I}}(x)$ to the top-$K$ experts selected from the combined pool of intra-vision and inter-modality groups $E^{\mathcal{I/S}}$, while textual tokens are routed by the textual router $G^{\mathcal{T}}(x)$ to the top-$K$ experts from the intra-language and inter-modality groups $E^{\mathcal{T/S}}$. The routing mechanism enables flexible expert selection patterns: each token may activate only intra-modality experts, only inter-modality experts, or a combination of both. This hierarchical routing strategy ensures efficient multi-modal learning while simultaneously modeling both modality-specific features and cross-modal associations. All experts in $\mathcal{E}$ are initialized from the LLM’s FFN layers, preserving prior knowledge while enabling adaptive multi-modal specialization.
2.3 Training Recipe
To balance simplicity and efficacy, we propose a two-stage training strategy that effectively activates the capabilities of the MoIIE architecture. In the first stage, we focus on pretraining the connection module to align visual representations with the LLM’s linguistic embedding space, enabling consistent processing of inputs across different modalities. In the second stage, we initialize both intra-modality and inter-modality experts using the pre-trained FFNs from the base LLM and fine-tune the entire model with diverse visual instruction datasets. During this stage, the training objectives include multi-modal expert learning for capturing modality-specific features and cross-modal associations, and optimizing versatile multi-modal understanding and instruction-following capability.
Training Objective
To ensure load balancing among experts within the MoIIE module while optimizing the model for overall multi-modal understanding, the training objective $\mathcal{L}$ is composed of two components: the language modeling cross-entropy loss $\mathcal{L}_{\text{lm}}$ and an auxiliary loss $\mathcal{L}_{\text{aux}}$ aimed at load balancing among experts MistralAITeam (2023).
$\mathcal{L}=\mathcal{L}_{\text{lm}}+\alpha\cdot\mathcal{L}_{\text{aux}},$ (3)
where $\alpha$ is a weighting coefficient set to 0.001. Notably, higher values of $\alpha$ adversely impact model performance, as detailed in Section 3.6. The auxiliary loss $\mathcal{L}_{\text{aux}}$ is defined as:
$\mathcal{L}_{\text{aux}}=|\mathcal{E}|\cdot\sum_{i=1}^{|\mathcal{E}|}\left(\mathbb{E}_{x}\left[G_{i}(x)\right]\cdot\mathbb{E}_{x}\left[\mathbf{1}_{i}(x)\right]\right),$ (4)
$G_{i}(x)$ is the routing probability of assigning token $x$ to expert $E_{i}$, and $\mathbf{1}_{i}(x)$ denotes the indicator function that equals 1 when expert $E_{i}$ is activated for token $x$ and 0 otherwise.
3 Experiments
3.1 Implementation Details
We employ a pre-trained SigLIP model Zhai et al. (2023b) and CLIP Radford et al. (2021)in Appendix A.4 as visual encoder, a two-layer MLP as connection module. For the LLM backbone, we utilize phi-3-mini Marah Abdin (2024) and LLaMA3-8b AI@Meta (2024). Our method and all compared MoE architectures use a 4-experts configuration. Specifically, our MoIIE module includes 2 intra-modality experts (one for vision and one for language) and 2 inter-modality experts. More detailed discussion of expert settings is provided in the ablation studies 3.6.
[…]
Data Backbone General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
Dense 2M Phi-3-mini 74.4 62.7 81.6 42.9 70.2 86.2 28.3 65.1 41.9 53.2 64.3 41.3 30.8 57.1
[…]
Dense 2.7M Phi-3-mini 75.0 63.4 81.6 41.1 70.1 87.0 31.7 64.6 48.2 57.5 73.7 40.1 31.0 58.8
Vanilla MoE 2.7M Phi-3-mini 75.2 63.3 81.4 42.2 70.6 87.2 30.8 65.2 47.8 57.2 74.7 42.2 31.2 59.1
Modality MoE 2.7M Phi-3-mini 73.9 63.5 81.7 40.1 70.4 87.1 32.9 65.7 48.9 58.8 74.1 40.4 31.3 59.1
[…]
Dense 2M LLaMA3-8B 74.8 64.8 82.2 47.9 72.1 86.6 30.1 67.2 43.4 54.0 76.6 41.1 30.9 59.4
Vanilla MoE 2M LLaMA3-8B 75.5 64.5 82.1 46.0 71.6 86.3 33.0 67.3 43.8 54.7 76.2 40.7 30.6 59.4
Modality MoE 2M LLaMA3-8B 75.6 64.3 82.3 46.3 72.3 86.7 35.0 67.6 44.9 55.4 76.7 41.6 31.2 60.0
[…]
During training, we utilize the AdamW optimizer Kingma and Ba (2017) with a cosine learning rate scheduler for one epoch across both stages. In the first stage, only the connection module is optimized with a learning rate of $1\times 10^{-3}$, using the Bunny-pretrain-LAION-2M dataset He et al. (2024). In the second stage, all parameters are unfrozen for joint SFT and sparse upcycling training. The learning rate is set to $2\times 10^{-6}$ for the visual encoder and $2\times 10^{-5}$ for all other components. For this stage, we attempt training with different scales of visual instruction data: (1) 1.3M samples from MGM-Instruction Li et al. (2024c); (2) 2M samples by adding Bunny-695K He et al. (2024); (3) 2.7M samples by further incorporating LLaVA-NEXT-779k Liu et al. (2024b)(4)8M advanced training data from LLaVA-OV Li et al. (2024a). Training is conducted using DeepSpeed Shuaiwen Leon Song (2023) with ZeRO-3 optimization. All compared baselines in the main results and our method follow the same training setup and data configurations. Further implementation details and results are provided in Appendix A.
3.2 Evaluation Benchmarks
We conduct a comprehensive evaluation across 13 diverse multi-modal benchmarks. Specifically, general multi-modal benchmarks include MMBench-EN Liu et al. (2024c), MM-Vet Yu et al. (2024), GQA Hudson and Manning (2019), VQAv2 and SEED-Image  Li et al. (2023a). For knowledge-based question answering, we utilize MMMU validation subset Yue et al. (2024), AI2D Kembhavi et al. (2016),SciQA-IMG Lu et al. (2022) and Mathvista Lu et al. (2024b). For OCR-based question answering, we assess performance on TextVQA Singh et al. (2019), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021). Additionally, we evaluate hallucination robustness using POPE Li et al. (2023c) and HallusionBench Guan et al. (2024).
3.3 Main Results
We compare our MoIIE-equipped model against three baseline variants: the original dense model (Dense) without MoE, the vanilla MoE model MistralAITeam (2023) that routes tokens from all modalities to a shared pool of experts (Vanilla MoE), and the model with modality-specific experts where text and image tokens are routed to their respective specialized experts (Modality MoE). The comparisons are conducted across multiple visual instruction tuning data scales and varying LLM backbones, as shown in Table 3.1. The key findings are as follows:
(1) The MoIIE module consistently outperforms all other architectural variants across most benchmarks with different data scales, visual encoders, and LLM backbone sizes, demonstrating the powerful generalization capability of our MoIIE. Integrating both modality-specific features and cross-modal associations can effectively enhance performance on diverse multi-modal tasks. Further analysis of expert modules is provided in Appendix A.3.
(2) Specifically, our MoIIE module is particularly effective on knowledge-based QA and hallucination benchmarks that requiring sophisticated cross-modal interaction between textual concepts and corresponding visual entity regions. In contrast, OCR-based QA tasks primarily involve interpreting information verbatim from images, where Modality MoE can already achieves competitive performance. This performance difference highlights MoIIE’s capacity to handle both modality-specific features and cross-modal associations.
(3) We further provide detailed performance trends with increasing fine-tuning data in Figure 3. As the data scale continues to increase, our MoIIE progressively improves while other architectures, especially Dense and Modality MoE, encounter performance limitations. Moreover, with larger training datasets, our MoIIE consistently outperforms all alternatives, with the performance gap between MoIIE and other architectures widening. This suggests that the MoIIE framework offers superior scaling properties for multi-modal learning, effectively leveraging larger datasets to enhance representation power without the parameter inefficiency of dense models or the limited cross-modal reasoning of strictly modality-separated experts.
Figure 3: Performance variation across different visual instruction tuning data scales . MoIIE outperforms other architectural variants in achieving superior scaling efficiency.
3.4 Comparison with MoE-based LLMs Initialized LVLMs
We compare our MoIIE, pre-trained on the LLaVA-OV-Single-Image dataset Li et al. (2024a), with open-source LVLMs initialized from MoE-based LLMs. Models such as SPHINX-X Liu et al. (2024a), MGM Li et al. (2024c), and CuMO Li et al. (2024b), built on Mixtral-8x7B MistralAITeam (2023), exploit large-scale language data for robust initialization but require substantial computational resources, limiting practicality in constrained settings. By contrast, MoIIE attains comparable performance across diverse multimodal benchmarks with 17% fewer activated parameters, thereby reducing both training and deployment costs. Moreover, it enables constructing MoE-based LVLMs from arbitrary LLM backbones, offering flexible expert configurations without performance degradation.
Table 2: Comparisons of MoIIE with open-source LVLMs utilizing pre-trained MoE-based LLM backbones. In addition to performance gains, our MoIIE offers the flexibility to be constructed using any dense LLM backbone. “Act. Param” indicates activated parameters while “All Param” indicates the total number of model parameters.
[…]
SPHINX-X Liu et al. (2024a) - $\sim$40B 68.0 63.8 81.1 89.6 71.3 40.9 31.1 42.7 74.5
MGM Li et al. (2024c) 13.5B $\sim$40B 69.2 - - - 75.6 45.8 41.8 41.8 -
CuMO Li et al. (2024b) 13.5B $\sim$40B 66.0 63.8 81.8 85.7 75.3 48.7 45.0 38.2 77.9
[…]
3.5 The Discussion of Training Pipeline
Converting dense LLMs into MoE-based LVLMs typically follows a three-stage pipeline Lin et al. (2024a); Chen et al. (2024b): (1) multimodal pretraining, (2) supervised fine-tuning (SFT) with multimodal instructions, and (3) sparse upcycling to obtain MoE-based LVLMs. This process is complex, and limited training in the final stage often weakens generalization. We propose a simplified two-stage approach: after pretraining, SFT and sparsification are combined into one phase, enabling direct activation of both multimodal and MoE features. Specifically, the three-stage pipeline uses 2M visual instruction samples for SFT (tuning all parameters) and LLaVA-NEXT-779k for sparse upcycling (only tuning MoE layers), while our two-stage method uses 2.7M samples for joint SFT and sparse upcycling (tuning all parameters). As shown in Table 3, our approach consistently outperforms the three-stage pipeline under the same conditions, offering both simplicity and improved effectiveness.
Table 3: Comparisons of different training strategies. Pretrain-SFT-MoE represents the three-stage training method where MoE denotes sparse upcycling from multi-modal dense checkpoints, tuning only MoE layers Lin et al. (2024a) after SFT. Pretrain-SFT&MoE is our proposed two-stage training method, in which sparse upcycling is integrated with SFT, tuning the full model parameters simultaneously.
[…]
Method Stage General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
Pretrain-SFT-MoE 3 75.6 63.1 81.6 41.2 70.3 86.9 31.0 64.8 48.2 56.2 63.8 39.3 31.3 57.9
[…]
3.6 Ablation Study
Table 4: Ablation study for various MoIIE module configurations. This table compares the performance of different MoIIE setups, including variations in the MoIIE location, load balance coefficient, number of experts and experts settings.
[…]
Number of Experts 4 8 65.1 75.6 63.4 42.2 86.5 43.0 71.2 63.9(+0.2)
Experts Balance(8 experts) Balance Unbalance 65.4 75.0 63.4 41.1 86.8 41.3 71.6 63.5(-0.2)
MoIIE Module Location Interleaved Full 65.8 75.8 63.1 40.3 87.0 42.6 70.6 63.6(-0.1)
Load Balance Coefficient 0.001 0.01 65.5 75.4 63.2 39.0 86.4 42.0 70.7 63.2(-0.5)
[…]
The Impact of the Number of Experts
Increasing the number of experts generally enhances MoE performance Lepikhin et al. (2020); FedusF et al. (2022). As shown in Table 4, expanding the MoIIE expert pool from 4 to 8 yields gains of 0.3, 0.2, and 0.9 points on MMBench, GQA, and MMMU, respectively, while results on POPE and TextVQA remain stable or slightly decline. These findings suggest that larger expert pools improve generalization on complex benchmarks but may require more data to fully realize their capacity.
The Impact of Experts Balance
Maintaining a balanced allocation of intra- and inter-modality experts is crucial for performance. As shown in Table 4, under 8-expert setup, Balanced configuration (2 vision, 2 language, 4 shared; 1:1 ratio) consistently outperforms the Unbalanced configuration (3 vision, 3 language, 2 shared; 3:1 ratio). The balanced design ensures more effective cross-modal interaction, getting stronger overall performance.
Layer Location of MoIIE module
We investigate two configurations for integrating MoIIE into LVLMs: an interleaved design, where MoIIE modules are inserted every other layer alongside dense layers, and a full design, where all layers are replaced with MoIIE modules Lin et al. (2024a); AI@Meta (2025). As shown in Table 4, the full design does not yield notable performance gains but significantly increases training cost and computational overhead due to the larger parameter count. By contrast, the interleaved configuration achieves comparable results with substantially higher efficiency and lower resource usage. We therefore adopt the interleaved design as the optimal balance of performance, efficiency, and scalability for MoIIE-based LVLMs.
Impact of Load Balance Coefficient
In MoE architectures, there is a trade-off between load balancing and performance. While the auxiliary loss promotes balanced expert usage Cai et al. (2024), excessive balancing can hinder multimodal learning and weaken cross-modal associations. As shown in Table 4, increasing the auxiliary loss weight $\alpha$ consistently reduces performance, especially on MMMUval and MM-Vet. These results underscore the importance of tuning $\alpha$ to maintain expert specialization and robust multimodal understanding.
3.7 Visualization of Token Pathways
Figure 4: Visualization of experts activated pathways. The figure shows the top-2 activated experts for text and image tokens, with Expert V and Expert L are intra-modality experts, Expert S are inter-modality experts.
As shown in Figure 4, we visualize expert activation pathways for image and text tokens in the MoIIE module on the MME test set. Distinct layer-wise patterns emerge: in shallow layers, tokens primarily activate intra-modality experts (Expert V for vision, Expert L for text), reflecting a focus on modality-specific feature extraction with limited cross-modal interaction. In deeper layers, activation shifts toward inter-modality experts (Expert S), enabling stronger cross-modal fusion. Notably, both modalities often converge on the same inter-modality experts (e.g., Expert S_1 in the 9th layer, Expert S_2 in the 12th). These results demonstrate that MoIIE dynamically balances modality-specific and cross-modal processing, supporting early-stage representation learning and late-stage integration.
4 Related Works
Large Vision-Language Models
Recent advances in LLMs Bai et al. (2023a); AI@Meta (2024); Touvron et al. (2023b); Chiang et al. (2023); Jiang et al. (2023); Bi et al. (2024); Team (2023); OpenAI (2023b, a); Touvron et al. (2023a) have shown strong generalization and instruction-following abilities, spurring research on Large Vision-Language Models (LVLMs) Li et al. (2023b); Alayrac et al. (2022); Liu et al. (2023a, b, 2024b). Progress in LVLMs has been driven by high-quality data He et al. (2024); Chen et al. (2024a); Zhang et al. (2024), extended training schemes Bai et al. (2023b); Lu et al. (2024a); Laurençon et al. (2024), support for high-resolution inputs Li et al. (2024e); Wang et al. (2024b), and multi-encoder designs Shi et al. (2024); Liu et al. (2024a); Fan et al. (2024). The latest open-source LVLMs Li et al. (2024a); Wang et al. (2024a); Chen et al. (2025) achieve state-of-the-art results by integrating large-scale datasets, dynamic resolution, and advanced LLMs.
Mixture of Experts in LLMs
The MoE paradigm Jacobs et al. (1991) scales model capacity by activating only a subset of experts, balancing efficiency and performance. It typically replaces feed-forward layers with expert modules and relies on Top-K routing Lepikhin et al. (2020); Du et al. (2022); FedusF et al. (2022); Zoph et al. (2022); Rajbhandari et al. (2022); Xue et al. (2024). Sparse upcycling Komatsuzaki et al. (2023) further reduces costs by converting dense models into sparse ones. Recent MoE-based LLMs xAI (2024); MistralAITeam (2023); Muennighoff et al. (2024); DeepSeek-AI (2024a, b) improve stability and efficiency through large-scale training strategies.
Mixture of Experts in LVLMs
MoE has also been applied to LVLMs, e.g., ARIA Li et al. (2025) and DeepSeek-VL2 Wu et al. (2024), which leverage MoE-based LLM backbones MistralAITeam (2023); DeepSeek-AI (2024a) but remain constrained by fixed expert settings and high training costs. Sparse upcycling Chen et al. (2024b); Lin et al. (2024a, c) improves scalability, yet struggles to capture both modality-specific and cross-modal interactions. More recent work Zhou et al. (2025) extends parameter-efficient fine-tuning to multimodal experts, but the role of modality-specific expert groups remains underexplored. To address these gaps, we propose a modality-aware MoE architecture that integrates intra- and inter-modality experts. Combined with a two-stage training strategy, our approach enables flexible expert allocation and more scalable, high-performance LVLMs.
5 Conclusion
In this study, we propose MoIIE, a novel Mixture-of-Experts architecture paired with a simple yet effective two-stage training strategy that enables the flexible construction of powerful MoE-based LVLMs from any dense LLM. MoIIE organizes experts into specialized intra-modality and shared inter-modality groups to create dedicated pathways for both visual and linguistic processing while enabling rich cross-modal associations, significantly enhancing performance across a wide range of multi-modal tasks. Extensive experiments demonstrate that our framework achieves superior scaling efficiency and performance compared to existing dense and sparse LVLM architectures.
Limitations
A primary limitation of our work lies in the relatively limited training data and its exclusive focus on vision-language modalities. In future work, we plan to address these limitations by incorporating more high-quality multi-task datasets, supporting dynamic input resolutions, and extending MoIIE to broader modalities such as speech, thereby advancing its generalization and applicability in real-world multi-modal scenarios.
Appendix A Appendix
Figure 5: Comparison between the Vanilla MoE and our MoIIE. Left: The vanilla MoE module MistralAITeam (2023) routes all modality tokens into a single group of experts. Right: The MoIIE module introduces intra-modality and inter-modality experts group. Intra-modality Expert V for image tokens and Expert L for text tokens modeling modality specfic features. Inter-modality experts (Expert S) that process tokens from both modalities, modeling cross-modal associations, capable of handling tokens from both modalities. The visual router routes image tokens to Expert V and Expert S, while the textual router routes text tokens to Expert L and Expert S.
A.1 Visualization of activated experts
As illustrated in the figure6, we visualize the proportion of activated experts across four representative multimodal understanding benchmarks. For general QA, we use MME Fu et al. (2024); for OCR-based QA, TextVQA Singh et al. (2019); for visual-centric reasoning, we use MMVP Tong et al. (2024); and for hallucination, we use HallusionBench Guan et al. (2024).
We consistently observe that lower layers of the model tend to activate modality-specific visual experts, indicating specialized modeling of image features. In contrast, as the layer depth increases, the activation ratio of modality-specific experts decreases, while the use of modality-shared experts increases, suggesting a shift toward modeling shared semantic information across modalities.
Figure 6: Visualization of the proportion of activated experts across diverse dimensions in multimodal understanding benchmarks.
A.2 Training Details
Our detailed training settings and hyper-parameters of MoIIE are shown in Table 5.
Configuration Stage 1 Stage 2
Experts type - FFN
Experts -
1 visual expert
2 shared experts
1 textual expert
Top-K - 2
Visual encoder siglip-so400m-patch14-384
Connection module 2 Linear layers with GeLU
Image resolution 384 x 384
Learning rate 1e-3 {connection module}
2e-5 {LLM,connection module}
2e-6 {visual encoder}
LR schedule Cosine decay
Weight decay 0
Optimizer AdamW
Warmup ratio 0.03
Epoch 1
Global batch size 256 128
Deepspeed Zero2 Zero3
Max token length 4096
Table 5: Detailed training hyperparameters of MoIIE.
A.3 Expert Analysis
MMBench TextVQA MMMU_val
Textual experts 72.9 63.1 41.5
Visual experts 73.6 64 41
Shared experts 74.1 63.1 41.3
Table 6: Expert analysis between modality-specific experts and modalitt-shared experts.
To evaluate the effectiveness of modality-specific and modality-shared experts, we extracted checkpoints from the shared experts in MoIIE, as well as the modality-specific visual and textual experts from the Modality Expert model. These were tested on the MMbench Liu et al. (2024c) and TextVQA Singh et al. (2019) and MMMU validation subset Yue et al. (2024) which for General, OCR and konwledge based abilities of model.
As shown in Table 6 We find that modality-shared experts outperform modality-specific experts on general multi-modal benchmarks like MMbench Liu et al. (2024c), which require cross-modal reasoning. In contrast, visual experts perform better on OCR-based QA like TextVQA Singh et al. (2019), which emphasizes image details and modality-specific information whether texual experts are capable of knowledge-based QA like MMMU validation subset Yue et al. (2024), which requires the original textual knowledge of model. This further validates the effectiveness of our MoIIE architecture, which demonstrates that successful performance across a wide range of multi-modal tasks requires modeling both cross-modal associations and modality-specific features.
[…]
Data Visual Encoder General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
[…]
Dense 2.7M CLIP-L-336 74.5 63.0 81.3 40.9 69.9 86.5 31.2 64.1 47.7 57.1 73.2 39.6 30.6 58.4
Vanilla MoE 2.7M CLIP-L-336 74.6 63.1 81.1 41.0 69.7 86.5 30.9 64.3 47.5 57.0 74.0 40.5 31.0 58.5
Modality MoE 2.7M CLIP-L-336 74.0 63.0 81.5 39.7 69.5 86.6 32.3 64.4 48.1 57.7 73.7 40.1 31.3 58.6
[…]
Dense 4M SigLIP-So400m-384 76.2 64.6 82.5 42.3 71.4 88.1 32.4 66.1 49.4 58.8 75.0 41.3 32.1 60.0
Vanilla MoE 4M SigLIP-So400m-384 76.4 64.9 82.9 42.9 71.6 88.9 33.8 66.2 50.2 59.7 75.8 43.0 33.3 60.7
Modality MoE 4M SigLIP-So400m-384 75.1 64.7 82.7 41.9 71.3 88.0 35.3 67.4 50.8 59.0 75.3 41.8 32.5 60.4
[…]
A.4 More Experimental Results
we conduct extended ablation experiments using a different vision encoder(CLIP-L-336) as shown in Table A.3, replacing the vision encoder does not affect our main conclusions. Under identical training conditions,the proposed MoIIE architecture consistently outperforms all other ablation variants across most of the 13 benchmarks and shows a clear advantage in terms of average performance.
We also adopt a larger dataset that matches the scale of modern open multimodal benchmarks. Specifically, we followed LLaVA-OV Li et al. (2024a), which provides the high overall data quality among these open datasets a total training corpus of approximately 8M samples, 4M for insturct tuing stage, and used it as our expanded training corpus.
As shown in Table A.3, using a higher-quality and larger-scale dataset but still identical training conditions across all variants, our proposed MoIIE architecture continues to outperform all other variants across all the 13 multimodal benchmarks. Moreover, MoIIE achieves a more pronounced improvement in average performance (+1.2%) relative to the other variants. with the trainig data more advanced, the performance gap between MoIIE and other architectures widening.
These results demonstrate that MoIIE not only maintains its advantages under stronger training condition, but also scales more effectively in multimodal training, confirming its robustness and generality.